OpenStudy (anonymous):

Breck Hardware purchased from Black and Decker 10 Dust Busters for $24.95 each. What should Breck charge its customers for each Dust Buster if it has a 39.5% markup on cost.

OpenStudy (anonymous):

24.95/10=2.495 2.495*1.395=3.480525 A dust buster would cost $3.48

OpenStudy (anonymous):

oh whoops they charged for 24.95 for each

OpenStudy (anonymous):

So it should just be 24.95*1.395=34.48 Breck should charge $34.48 per dust buster